PrettyEyes:

Rachel enjoys exercising outdoors. Today she walked miles in hours. What is Rachel’s unit walking rate in miles per hour and in hours per mile?

PrettyEyes:

the fractions are 5 2/3 2 2/3

Convert:

5 2/3 / 2 2/3 = 17/3 17/3 / 8/3 = 17/3 x 3/8 (reciprocal and flip sign) = 17/8 17/8 = 2 1/8 miles per hour and 8/17 hours per mile. (flip over fraction to find hours per mile)

PrettyEyes:

thank you!!!

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @PrettyEyes thank you!!! \(\color{#0cbb34}{\text{End of Quote}}\) I'm happy to help! :)